The length of a rectangle is 4 in. more than twice its width. If the perimeter of the rectangle is 56 in., find the width of the rectangle.
Answer: The width is ______

Answer by ptaylor(2198)

Let x=width
Then 2x+4=length
Perimeter of a rectangle =2W+2L. So:
2x+2(2x+4)=56 get rid of parens
2x+4x+8=56 subtract 8 from both sides
2x+4x+8-8=56-8 collect like terms
6x=48 divide both sides by 6
x=8 in--------------------------------width
2x+4=2*8+4=20 in--------------------length
CK
Perimeter equals 56 in
2(8)+2(20)=56
16+40=56
56=56

Hope this helps----ptaylor